OpenStudy (anonymous):

simplify 2x-(3*0.6+2.7) a. -2.5 b.2.5 C.0.2x+2.7 d. 0.2x-2.7

OpenStudy (texaschic101):

2x - (3 * 0.6 + 2.7) = 2x - (1.8 + 2.7) = 2x - (4.5) = 2x - 4.5 I am not getting any of your answer choices